OpenStudy (anonymous):

f(x)=x^2+1; g(x)=sqrtx-9 Find domain

OpenStudy (anonymous):

Domain of f(x): All real numbers. Domain of g(x): All real numbers >=9.

OpenStudy (anonymous):

Was -9 under the square root symbol too?

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

The domain of \(f\) is \(\mathbb R\). The domain of \(g\) is \(\{x\in\mathbb R:x\ge9\}\). This is assuming that \(g(x)=\sqrt{(x-9)}\) as opposed to \(g(x)=\sqrt x - 9\).

OpenStudy (anonymous):

If it's under the square root symbol then @myko is correct. If not, then my response is correct.

OpenStudy (anonymous):

Actually, your response is not correct in either case. Being a whole number has nothing to do with it. If it's not under the square root the answer is the positive reals.

OpenStudy (anonymous):

I advise you to use parentheses when possible.

OpenStudy (anonymous):

@nbouscal good point. I was thinking positive reals and wrote whole number

OpenStudy (anonymous):

Domain of f(x): All real numbers. Domain of g(x): All real numbers >=9. (If sqrt(x-9)) Domain of f(x): All real numbers. Domain of g(x): 0 and all positive real numbers (If sqrt(x) -9)